# Differential Communication in Channels with Mobility and Delay Spread using Zak-OTFS

### v1

@@ III Differential Communication in Zak-OTFS
=$\overset{\small(a)}{=}\frac{1}{MN}\sum_{k,l}\mathbf{h}_{\mathrm{eff}}[k,l]\sum_{k_{0},l_{0}}\mathbf{X}[k_{0},l_{0}]\times$
@@
where step ${\small(a)}$ follows because the pulsone samples and noise samples are uncorrelated. Substituting $\bar{n}=n-k$ in (11):
@@ IV-A Error Propagation
Fig. 3 plots the instantaneous NMSE of the channel estimate obtained using the DO frame (using 4-QAM symbols) at the receiver at low and high SNR values. From the plot, it is observed that at low SNR (SNR = 0 dB) error propagation results in instantaneous NMSE increasing with the number of frames. It is also observed that every 30 DO frames, there is a dip in the NMSE value, wherein the estimate from the pilot frame curbs error propagation. However, at high SNR (SNR = 25 dB), the number of errors is so low that there is no significant error propagation between two pilot transmissions.
@@
IV-B 4-QAM

### v2

@@ III Differential Communication in Zak-OTFS
=$\overset{(a)}{=}\frac{1}{MN}\sum_{k,l}\mathbf{h}_{\mathrm{eff}}[k,l]\sum_{k_{0},l_{0}}\mathbf{X}[k_{0},l_{0}]\times$
@@
where step $(a)$ follows because the pulsone samples and noise samples are uncorrelated. Substituting $\bar{n}=n-k$ in (11):
@@ IV Numerical Results
(a) Actual $\mathbf{h}_{\mathrm{eff}},M=11,N=13$
(b) Estimated $\mathbf{h}_{\mathrm{eff}}$ using DO with 4-QAM
(c) Estimated $\mathbf{h}_{\mathrm{eff}}$ using DO with 256-QAM
(d) Actual $\mathbf{h}_{\mathrm{eff}},M=31,N=37$
(e) Estimated $\mathbf{h}_{\mathrm{eff}}$ using DO with 4-QAM
(f) Estimated $\mathbf{h}_{\mathrm{eff}}$ using DO with 256-QAM
(g) Actual $\mathbf{h}_{\mathrm{eff}},M=101,N=103$
(h) Estimated $\mathbf{h}_{\mathrm{eff}}$ using DO with 4-QAM
(i) Estimated $\mathbf{h}_{\mathrm{eff}}$ using DO with 256-QAM
Figure 3: Demonstrating the effect of size of frame and constellation on the performance of the proposed differential communication scheme for data SNR of 25 dB. RRC pulse shaping with parameter $\beta_{\tau}=\beta_{\nu}=0.6$. Three frame sizes are considered $M=11,N=13$ (first row), $M=31,N=37$ (second row), and $M=101,N=103$ (third row). For the DO frame, constellation sizes of 4-QAM and 256-QAM are considered. No significant performance difference between 4-QAM and 256-QAM across frame sizes. Channel estimation accuracy improves with increasing frame size.
@@ IV-A Error Propagation
Figure 2 plots the instantaneous NMSE of the channel estimate obtained using the DO frame (using 4-QAM symbols) at the receiver at low and high SNR values. From the plot, it is observed that at low SNR (SNR = 0 dB) error propagation results in instantaneous NMSE increasing with the number of frames. It is also observed that every 30 frames, there is a dip in the NMSE value, wherein the estimate from the pilot frame curbs error propagation. However, at high SNR (SNR = 25 dB), the number of errors is low and there is no significant error propagation between two pilot transmissions. Note that, at SNR = 0 dB, the target NMSE required for performing close to perfect CSI is also very high, while at SNR = 25 dB, the target NMSE is lower to achieve close to perfect CSI performance and this trend is supported by the proposed receiver.
(a) NMSE
(b) BER
@@
IV-B Effect of Frame and Constellation Size on the Estimation Performance
Figure 3 plots the surface plot of the perfect channel and channel estimated using the DO frame for different frame sizes and constellation sizes. We consider a small frame with $M=11,N=13$, a medium frame with $M=31,N=37$, and a large frame with $M=101,N=103$. For each frame size, the channel estimated using the DO frame using 4-QAM (small constellation size) and 256-QAM (large constellation size) is plotted. It is seen that, for all frame sizes, the estimated channel does not vary significantly with change in constellation size but there is significant improvement with increase in frame size. This corroborates (16), wherein the estimate of the channel converges to the true channel asymptotically with frame size.
IV-C 4-QAM